A 3000-km long cable consists of 7 turns, with diameter of 0.73mm each. Calculate the resistance R of the cable. The resistivity of the copper wire is 3x10-6 Ω.cm.             


1
Expert's answer
2022-02-07T15:28:43-0500

The resistance is


"R=\dfrac{4n\rho L}{\pi d^2}=150\text M\Omega."